OpenStudy (anonymous):

Will medal and fan What is the sum of the geometric sequence -4, 24, -144 . . . if there are 6 terms? -159,963 26,660 26,661 159,963

OpenStudy (anonymous):

do 144 times neg 6 and that times neg 6 and that again times neg 6 to get all 6 terms

OpenStudy (anonymous):

then add them

OpenStudy (anonymous):

I did and the answer I got isn't the same as the options

OpenStudy (xapproachesinfinity):

you mean the sum of geometric series not sequence?

OpenStudy (anonymous):

No it says sequence

OpenStudy (xapproachesinfinity):

first you need to find all the 6 terms then sum up

OpenStudy (xapproachesinfinity):

you have three terms so you are missing three more to make 6 terms

OpenStudy (xapproachesinfinity):

what is the fourth term?

OpenStudy (anonymous):

864, -5184, 31104 thee are the next three terms. At least that's what I got. If I add them up I don't get one of the answers they provide

OpenStudy (xapproachesinfinity):

hmm really then you must of made a mistake

OpenStudy (xapproachesinfinity):

how did you get the fourth term for example

OpenStudy (anonymous):

\[-144 \times -6 = 864\]

OpenStudy (xapproachesinfinity):

yes!

OpenStudy (xapproachesinfinity):

you should get the right answer i don't where you missed

OpenStudy (xapproachesinfinity):

let's try doing the fomula \[S_n=\frac{a(1-r^n)}{(1-r)}\]

OpenStudy (xapproachesinfinity):

a=first term n=6 r=-6

OpenStudy (xapproachesinfinity):

@Tarynrachelle where you going?

OpenStudy (anonymous):

I'm still here I'm doing the problem

OpenStudy (xapproachesinfinity):

well low and behold i get one of the answer choices

OpenStudy (xapproachesinfinity):

you should get the right answer once you do the calculations

OpenStudy (anonymous):

The third answer?

OpenStudy (xapproachesinfinity):

and actually i get the correct answer even the method i told you in the start without formula. all of them should work since the formula came from it to begin with

OpenStudy (xapproachesinfinity):

no

OpenStudy (anonymous):

Idk what I'm doing wrong then.

OpenStudy (xapproachesinfinity):

well to me they all give the same answer, something must be wrong in yur work

OpenStudy (xapproachesinfinity):

anyway you got the formula, you got the other way just see where you had missed things

OpenStudy (xapproachesinfinity):

did you do it yet?

OpenStudy (xapproachesinfinity):

use the formula

OpenStudy (anonymous):

Okay well this is what I plug into the formula\[S _{n}=\frac{ 4(1-(-6)^{6}) }{ (1-(-6)) }\] am I wrong?

OpenStudy (xapproachesinfinity):

yeah

OpenStudy (anonymous):

That's supposed to be a -4

OpenStudy (xapproachesinfinity):

yes -4 not 4 just about to say that

OpenStudy (anonymous):

Is it still wrong even with the -4? Or is it correct with the -4?

OpenStudy (xapproachesinfinity):

it is correct with -4

OpenStudy (anonymous):

Okay. Hold on I think I see where I messed up

OpenStudy (anonymous):

Nope still not getting one of the answer choices..

OpenStudy (anonymous):

I plug it into my calculator and get 186614. That's the same answer I get when I add them all up

OpenStudy (anonymous):

Could you explain the steps for me so I know what I'm doing wrong?

OpenStudy (xapproachesinfinity):

hold a sec

OpenStudy (xapproachesinfinity):

ok back

OpenStudy (xapproachesinfinity):

how did you put it in your calc?

OpenStudy (anonymous):

\[-4(1-(6)^6)\div(1-(6))\]

OpenStudy (xapproachesinfinity):

no bottom is (1-(-6))=1+6

OpenStudy (xapproachesinfinity):

where are the negatives of 6?

OpenStudy (anonymous):

My calculator wasn't getting the right answer but this is it http://www.wolframalpha.com/input/?i=-4%281-%28-6%29%5E6%29%2F%281-%28-6%29%29

OpenStudy (xapproachesinfinity):

well for the top you didn;t have to worry a lot of that negative since you have an even power but bottom no

OpenStudy (anonymous):

It's the second choice

OpenStudy (xapproachesinfinity):

yes!

OpenStudy (xapproachesinfinity):

i think you put the wrong format into your calc

OpenStudy (anonymous):

Probably. But thank you

OpenStudy (xapproachesinfinity):

No problem